# Features are fate: a theory of transfer learning in high-dimensional regression

### v1

@@ 3.3 Linear transfer
We prove Theorem 3.7 in Appendix B.6. The structure of the result in Theorem 3.7 merits some discussion. After pretraining in the feature learning regime $\alpha\to 0$, the feature space of the network has sparsified so that it can only express functions along $\bm{\beta}_{\mathrm{s}}$ (Theorem 3.4). Since the features of the network cannot change in linear transfer, the main contribution to the generalization error is $\sin^{2}\theta$, which can be viewed as the norm of the projection of the target function into the space orthogonal to the features spanned by the pretrained network. This is an irreducible error that is the best case risk given that the features cannot change. The second term comes from the finiteness of the training set. Since linear transfer learns from a finite sample of training points, minimizing the training error can effectively “overfit the noise” and the learned function distorts away from the ground truth.
Luckily, since the pretrained feature space has sparsified, the effect of finite sampling and additive label noise decays as $\sim 1/n$, effectively filtering out the $d$-dimensional noise by projecting it onto a single vector. Compare this to the generalization of the scratch trained network (10). There, the features of the equivalent linear regression problem, $\bm{X}$, have support over all $d$-dimensions, so there is no irreducible error term. The expressivity, however, comes at a cost. Each dimension of the regression vector is vulnerable to noise in the training data, and the projection of the target function onto the feature space is strongly distorted due to finite sampling (i.e. $\sim\gamma$). We can precisely analyze this trade off by comparing (10) and (12). In the limit $n,d\to\infty$, the transferability (7) is
@@
which is plotted in Fig. 1(a). From (13) we can identify the regions of negative transfer for this model, which are shaded in red in Fig. 1(b). In the underparameterized regime ($\gamma>1$), there is negative transfer for all $\gamma-1>\frac{\sigma^{2}}{\sin^{2}\theta}$. In words, at fixed $\gamma$ and $\sigma$, i.e., fixing the number of data points and label noise, as the norm of the out-of-subspace component increases, transfer efficiency degrades.
In the overparameterized regime ($\gamma<1$), negative transfer only occurs when $\sigma<1$.This can be viewed as a condition on the signal-to-noise ratio of the target data: $\textrm{SNR}=\lVert\bm{\beta}_{\mathrm{t}}\rVert_{2}^{2}/\sigma^{2}=1/\sigma^{2}$. When $\textrm{SNR}<1$, scratch training can never recover the underlying vector $\bm{\beta}_{\mathrm{t}}$ and pretraining is always beneficial. When $\textrm{SNR}>1$, negative transfer occurs when $\theta\in(\arccos(1-\sigma),\pi/2)$ and $\gamma\in(\gamma_{+},\gamma_{-})$ where $\gamma_{\pm}=\frac{1}{2}[(1+\cos^{2}\theta-\sigma^{2})\pm\sqrt{(1+\cos^{2}\theta-\sigma^{2})^{2}-4\cos^{2}\theta}]$. In the noiseless case $\sigma\to 0$, this expression simplifies to $\theta\in(0,\pi/2)$, $\cos^{2}\theta<\gamma<1$ (see Appendix E Fig. 6). This condition requires that there is more data than the there is target function power in the direction learned during pretraining. As $\sigma$ increases, the region of negative transfer shrinks, since the noise corrupts the scratch trained accuracy. Finally we mention that the two regions of negative transfer in Fig. 1 are separated by positive transfer that persists even when $\theta=\pi/2$. We dub this effect anomalous positive transfer, since the pretrained features are completely orthogonal to those in the target, yet transferability is still positive. This effect stems from the double descent phenomenon in scratch training— even though linear transfer performance is as bad as possible, transferability is still positive, since the scratch trained model has infinite generalization error. In Appendix E Fig. 5 we demonstrate that the dataset based discrepency measures of Section 2.1 are indeed misleading: neither $D_{\mathrm{KL}}$, nor $W_{1}$ are negatively correlated with increased transferability.
@@ 3.3.1 Ridge regularization cannot fix negative transfer
In the following theorem, which we prove in Appendix B.7, we show that the generalization error for regularized linear transfer is a strictly increasing function of the ridge parameter $\lambda$, leading to a larger region of negative transfer for any $\lambda>0$ (Fig. 2).
@@ 3.4 Fine-tuning
Theorem 3.9 is proven in Appendix B.8. Theorem 3.9 yields an expression for the fine-tuning transferability, which is plotted in Fig. 3(a):
@@
When the model is underparameterized $\gamma>1$, there is a unique global minimum in the space of $\bm{\beta}=\bm{W}_{1}\bm{W}_{2}\cdots\bm{W}_{L}$. Since gradient flow converges to a global minimum, (Theorem 3.5), fine tuning loses the memory of the pretrained initialization leading to zero transferability (white region in Fig. 3(b)). When the network is overparameterized, however, there is a subspace of global minima. We show in the Section B.8 that the pretrained initialization induces an implicit bias of gradient flow away from the minimum norm solution. For $\theta<\pi/3$, the pretrained features are beneficial, leading to positive transfer. For $\theta>\pi/3$, however, the pretrained features bias the network too strongly toward the source task, leading to negative transfer.
@@ 4 Student-teacher ReLU networks
Figure 4: Linear transfer in two-layer ReLU networks We train a two layer ReLU network (18) with $m=1000$ neurons on a teacher with $m_{*}=100$ neurons and $d=100$ dimensional gaussian data, according to the ablated transfer setup (19), (20). For these experiments, we set the label noise $\sigma=0$. (a) Gram matrix from the kernel of the pretrained model (b) Gram matrix from the kernel of the ground truth source function $f^{*}_{\textrm{s}}(\bm{x})$. The two gram matrices are nearly indistinguishable suggesting that the kernel sparsifies to the represent features in the source task. (c) Generalization error of the scratch trained model as a function of dataset size $n$, fit to a power law (d) Norm of out-of-RKHS component of target function $\lVert P^{\perp}f^{*}_{\textrm{t}}(\bm{x})\rVert_{L_{2}}^{2}$, normalized by target function norm $\lVert f^{*}_{\textrm{t}}(\bm{x})\rVert_{L_{2}}^{2}$ as a function of excess target features $\mu$. (e) Heat map of transferability as a function of excess target features $\mu$ and dataset size $n$. We normalize the transferability by variance in the target data. Gray circles represent the point of negative transfer predicted by our theory. Results are averaged over $100$ realizations of the data and $10$ realizations of random draws of the teacher.
In the following, we demonstrate that many of the results from our analytically solvable model also hold, qualitatively, in the more complicated setting of linear transfer with nonlinear networks. In particular, we choose a model of the form
$f(\bm{x})=\frac{1}{m}\sum_{i=1}^{m}c_{i}\sigma(\bm{w}_{i}^{T}\bm{x})$ (18)
where $\sigma(y)=\max\{0,y\}$ is the ReLU activation. We scale the model by $1/m$ to place the network in the mean field, feature learning regime (Chizat et al., 2019; Mei et al., 2018; Rotskoff & Vanden-Eijnden, 2022). As in the deep linear model, we choose source and target functions that are representable by the network. That is, we study this model in the student teacher setting. To vary the level of feature space overlap between the source and target functions, we define a network of $m_{*}$ neurons for the target task, and generate the source network by ablating a fraction $\mu$ of the hidden neurons form the target. More precisely, let $\mathcal{A}$ be a uniformly random subset of the index set $\{1,2,\cdots m_{*}\}$ with $|\mathcal{A}|=\mu m_{*}$. Then
@@
Previous work (Rotskoff & Vanden-Eijnden, 2022; Mei et al., 2018; Chizat, 2020) has shown that in the overparameterized setting $m\gg m_{*}$, gradient flow will converge to a global minimizer of the population loss, so that $\lim_{m\to\infty}\lim_{t\to\infty}f(\bm{x})=f_{\textrm{s}}^{*}(\bm{x})$, which establishes that the trained network builds a representation of $f_{\textrm{s}}^{*}(\bm{x})$ in the mean field limit. This does not necessarily mean that all of the hidden neurons of the model converge to those of the teacher, since any superfluous weight directions can be eliminated by setting the corresponding output weight to zero. However, we demonstrate empirically in Fig.  4(a)-(b) that this relationship is preserved at the level of the model’s kernel, so that $k(\bm{x},\bm{x}^{\prime})=\frac{1}{m}\sum_{i=1}^{m}\sigma(\bm{w}^{T}_{i}\bm{x})\sigma(\bm{w}^{T}_{i}\bm{x}^{\prime})\approx\frac{1}{(1-\mu)m_{*}}\sum_{i\in\mathcal{A}^{c}}\sigma(\bm{w}^{*T}_{i}\bm{x})\sigma(\bm{w}^{*T}_{i}\bm{x}^{\prime})$. This observation is analogous to Theorem 3.4: training in the feature learning regime causes the model’s features to sparsify to those present in the target function.
@@
As expected, the norm of this projection increases monotonically with $\mu$ as shown in Fig. 4(d). We show how to compute this projection in Appendix C. In the deep linear setting, $\lVert P^{\perp}f^{*}_{\textrm{t}}(\bm{x})\rVert_{L_{2}}^{2}=\sin^{2}\theta$, and $C(n)\sim 1/n$. While the asymptotic, typical generalization error of kernel regression has been studied in (Canatar et al., 2021), for the purposes of estimating the generalization error of the transferred model, we assume here that this generalization error is dominated by this irreducible term for the large $n$ target dataset sizes we consider, just as we showed for the deep linear model.
However, an expression for the generalization error of the scratch-trained model is also needed to derive the transferability. We are not aware of a theory of generalization error for infinite width nonlinear networks trained on a finite data in the mean field regime. Intriguingly, however, we demonstrate empirically (Fig. 4(c)) that the generalization error obeys a power law $\mathcal{R}_{\textrm{sc}}\sim An^{-\nu}$ with $\nu=1.18$. By setting our theoretically predicted generalization error of our transferred model $\lVert P^{\perp}f^{*}_{\textrm{t}}(\bm{x})\rVert_{L_{2}}^{2}$ equal to the empirically observed scaling law $An^{-\nu}$ for our scratch-trained model, we can approximately identify the point of negative transfer in $n$ for any given $\mu$ (gray circles in Fig. 4(e)). It is clear from Fig. 4(e) that this heuristic for finding the boundary between positive and negative transfer becomes more accurate as the number of target points becomes large, since the $n$-dependent component of the kernel regression generalization error goes to zero in this limit. The phase diagram in Fig. 4 for noiseless ReLU networks resembles the phase diagram for linear transfer with deep linear networks in the noiseless setting with $\sigma=0$ (Appendix E Fig.6. 7). Overall, this demonstrates that we are able to predict the phase boundary between positive and negative transfer in the ReLU case, using our conceptual understanding in the deep linear case.
@@ 5 Conclusion
In this paper, we highlight the importance of thinking about transfer learning in the context of the feature space of the pretrained model. In particular, we show that certain Integral Probability Metrics and $\phi$-divergence can be misleading when it comes to predicting transfer learning performance using the datasets alone. We then rigorously identify the number of data points necessary for transfer learning to outperform scratch training as a function of feature space overlap in deep linear networks. Finally, we demonstrate that our understanding of linear transfer carries over to shallow nonlinear networks as well.
Appendix A Initialization assumption
Following Yun et al. (2021) we place the following constraint on the initialization for some $\lambda>0$.

### v2

@@ 3.3 Linear transfer
We prove Theorem 3.7 in Appendix D.6. The structure of the result in Theorem 3.7 merits some discussion. After pretraining in the feature learning regime $\alpha\to 0$, the feature space of the network has sparsified so that it can only express functions along $\bm{\beta}_{\mathrm{s}}$ (Theorem 3.4). Since the features of the network cannot change in linear transfer, the main contribution to the generalization error is $\sin^{2}\theta$, which can be viewed as the norm of the projection of the target function into the space orthogonal to the features spanned by the pretrained network. This is an irreducible error that is the best case risk given that the features cannot change. The second term comes from the finiteness of the training set. Since linear transfer learns from a finite sample of training points, minimizing the training error can overfit the noise and the learned function distorts away from the ground truth. Luckily, since the pretrained feature space has sparsified, the effect of finite sampling and additive label noise decays as $\sim 1/n$, effectively filtering out the $d$-dimensional noise by projecting it onto a single vector. Compare this to the generalization of the scratch trained network (8). There, the features of the equivalent linear regression problem, $\bm{X}$, have support over all $d$-dimensions, so there is no irreducible error term. The expressivity, however, comes at a cost. Each dimension of the regression vector is vulnerable to noise in the training data, and the projection of the target function onto the feature space is strongly distorted due to finite sampling (i.e. $\sim\gamma$). We can precisely analyze this trade off by comparing (8) and (10). In the limit $n,d\to\infty$, the transferability (5) is
@@
which is plotted in Fig. 1(a). From (11) we can identify the regions of negative transfer for this model, which are shaded in red in Fig. 1(b). In the underparameterized regime ($\gamma>1$), there is negative transfer for all $\gamma-1>\frac{\sigma^{2}}{\sin^{2}\theta}$. In words, at fixed $\gamma$ and $\sigma$, i.e., fixing the number of data points and label noise, transfer efficiency degrades as the norm of the out-of-subspace component increases.
In the overparameterized regime ($\gamma<1$), negative transfer only occurs when $\sigma<1$.This can be viewed as a condition on the signal-to-noise ratio of the target data: $\textrm{SNR}=\lVert\bm{\beta}_{\mathrm{t}}\rVert_{2}^{2}/\sigma^{2}=1/\sigma^{2}$. When $\textrm{SNR}<1$, scratch training can never recover the underlying vector $\bm{\beta}_{\mathrm{t}}$ and pretraining is always beneficial. When $\textrm{SNR}>1$, negative transfer occurs when $\theta\in(\arccos(1-\sigma),\pi/2)$ and $\gamma\in(\gamma_{+},\gamma_{-})$ where $\gamma_{\pm}=\frac{1}{2}[(1+\cos^{2}\theta-\sigma^{2})\pm\sqrt{(1+\cos^{2}\theta-\sigma^{2})^{2}-4\cos^{2}\theta}]$. In the noiseless case $\sigma\to 0$, this expression simplifies to $\theta\in(0,\pi/2)$, $\cos^{2}\theta<\gamma<1$ (see Appendix G Fig. 7). We can view $\gamma$ as a dimensionless measure of the amount of target data, and $\cos^{2}\theta$ as the amount of power the target function has in the subspace of the pretraining task. The condition for negative transfer is satisfied when there is more target data than there is power in the pretrained subspace. As $\sigma$ increases, the region of negative transfer shrinks, since the noise corrupts the scratch trained accuracy. Finally we mention that the two regions of negative transfer in Fig. 1 are separated by positive transfer that persists even when $\theta=\pi/2$. We dub this effect anomalous positive transfer, since the pretrained features are completely orthogonal to those in the target, yet transferability is still positive. In this regime, transfer is positive soley because of the disproportionately large amount of data in the source task, not because pretraining learned useful features for the downstream task. By comparing the transferred model to a regularized scratch-trained model, we can eliminate this effect, which we show in simulation in Appendix G Fig. 5. In Appendix G Fig. 4 we demonstrate that distributional discrepency measures are indeed misleading: neither $D_{\mathrm{KL}}$, nor $W_{1}$ are negatively correlated with increased transferability.
@@ 3.3.1 Ridge regularization cannot fix negative transfer
In the following theorem, which we prove in Appendix D.7, we show that the generalization error for regularized linear transfer is a strictly increasing function of the ridge parameter $\lambda$, leading to a larger region of negative transfer for any $\lambda>0$ (Fig. 6).
@@ 3.4 Fine-tuning
Theorem 3.9 is proven in Appendix D.8. Theorem 3.9 yields an expression for the fine-tuning transferability, which is plotted in Fig. 2(a):
@@
When the model is underparameterized $\gamma>1$, there is a unique global minimum in the space of $\bm{\beta}=\bm{W}_{1}\bm{W}_{2}\cdots\bm{W}_{L}$. Since gradient flow converges to a global minimum, (Theorem 3.5), fine tuning loses the memory of the pretrained initialization leading to zero transferability (white region in Fig. 2(b)). When the network is overparameterized, however, there is a subspace of global minima. We show in the Section D.8 that the pretrained initialization induces an implicit bias of gradient flow away from the minimum norm solution. When the target vector has the majority of its norm along the pretraining subspace, i.e. $\cos\theta>1/2$, the pretrained features are beneficial, leading to positive transfer. For $\cos\theta<1/2$, however, the pretrained features bias the network too strongly toward the source task, leading to negative transfer.
@@ 4 Student-teacher ReLU networks
Figure 3: Linear transfer in two-layer ReLU networks We train a two layer ReLU network with $m=1000$ neurons on a teacher with $m_{*}=100$ neurons and $d=100$ dimensional gaussian data, according to the ablated transfer setup (16), (17). For these experiments, we set the label noise $\sigma=0$. (a) Gram matrix from the kernel of the pretrained model (b) Gram matrix from the kernel of the ground truth source function $f^{*}_{\textrm{s}}(\bm{x})$. The two gram matrices are nearly indistinguishable suggesting that the kernel sparsifies to the represent features in the source task. (c) Generalization error of the scratch trained model as a function of dataset size $n$, fit to a power law (d) Norm of out-of-RKHS component of target function $\lVert P^{\perp}f^{*}_{\textrm{t}}(\bm{x})\rVert_{L_{2}}^{2}$, normalized by target function norm $\lVert f^{*}_{\textrm{t}}(\bm{x})\rVert_{L_{2}}^{2}$ as a function of excess target features $\mu$. (e) Heat map of transferability as a function of excess target features $\mu$ and dataset size $n$. We normalize the transferability by variance in the target data. Gray circles represent the point of negative transfer predicted by our theory. Results are averaged over $100$ realizations of the data and $10$ realizations of random draws of the teacher.
In the following, we demonstrate that many of the results from our analytically solvable model also hold, qualitatively, in the more complicated setting of linear transfer with nonlinear networks. In particular, we choose a model of the form $f(\bm{x})=\frac{1}{m}\sum_{i=1}^{m}c_{i}\sigma(\bm{w}_{i}^{T}\bm{x})$ where $\sigma(y)=\max\{0,y\}$ is the ReLU activation. We scale the model by $1/m$ to place the network in the mean field, feature learning regime (Chizat et al., 2019; Mei et al., 2018; Rotskoff & Vanden-Eijnden, 2022). As in the deep linear model, we choose source and target functions that are representable by the network. That is, we study this model in the student teacher setting. To vary the level of feature space overlap between the source and target functions, we define a network of $m_{*}$ neurons for the target task, and generate the source network by ablating a fraction $\mu$ of the hidden neurons form the target. More precisely, let $\mathcal{A}$ be a uniformly random subset of the index set $\{1,2,\cdots m_{*}\}$ with $|\mathcal{A}|=\mu m_{*}$. Then
@@
Previous work (Rotskoff & Vanden-Eijnden, 2022; Mei et al., 2018; Chizat, 2020) has shown that in the overparameterized setting $m\gg m_{*}$, gradient flow will converge to a global minimizer of the population loss, so that $\lim_{m\to\infty}\lim_{t\to\infty}f(\bm{x})=f_{\textrm{s}}^{*}(\bm{x})$, which establishes that the trained network builds a representation of $f_{\textrm{s}}^{*}(\bm{x})$ in the mean field limit. This does not necessarily mean that all of the hidden neurons of the model converge to those of the teacher, since any superfluous weight directions can be eliminated by setting the corresponding output weight to zero. However, we demonstrate empirically in Fig.  3(a)-(b) that this relationship is preserved at the level of the model’s kernel, so that $k(\bm{x},\bm{x}^{\prime})=\frac{1}{m}\sum_{i=1}^{m}\sigma(\bm{w}^{T}_{i}\bm{x})\sigma(\bm{w}^{T}_{i}\bm{x}^{\prime})\approx\frac{1}{(1-\mu)m_{*}}\sum_{i\in\mathcal{A}^{c}}\sigma(\bm{w}^{*T}_{i}\bm{x})\sigma(\bm{w}^{*T}_{i}\bm{x}^{\prime})$. This observation is analogous to Theorem 3.4: training in the feature learning regime causes the model’s features to sparsify to those present in the target function.
@@
As expected, the norm of this projection increases monotonically with $\mu$ as shown in Fig. 3(d). We show how to compute this projection in Appendix E. In the deep linear setting, $\lVert P^{\perp}f^{*}_{\textrm{t}}(\bm{x})\rVert_{L_{2}}^{2}=\sin^{2}\theta$, and $C(n)\sim 1/n$. While the asymptotic, typical generalization error of kernel regression has been studied in (Canatar et al., 2021), for the purposes of estimating the generalization error of the transferred model, we assume here that this generalization error is dominated by this irreducible term for the large $n$ target dataset sizes we consider, just as we showed for the deep linear model.
However, an expression for the generalization error of the scratch-trained model is also needed to derive the transferability. We are not aware of a theory of generalization error for infinite width nonlinear networks trained on a finite data in the mean field regime. Intriguingly, however, we demonstrate empirically (Fig. 3(c)) that the generalization error obeys a power law $\mathcal{R}_{\textrm{sc}}\sim An^{-\nu}$ with $\nu=1.18$. By setting our theoretically predicted generalization error of our transferred model $\lVert P^{\perp}f^{*}_{\textrm{t}}(\bm{x})\rVert_{L_{2}}^{2}$ equal to the empirically observed scaling law $An^{-\nu}$ for our scratch-trained model, we can approximately identify the point of negative transfer in $n$ for any given $\mu$ (gray circles in Fig. 3(e)). It is clear from Fig. 3(e) that this heuristic for finding the boundary between positive and negative transfer becomes more accurate as the number of target points becomes large, since the $n$-dependent component of the kernel regression generalization error goes to zero in this limit. The phase diagram in Fig. 3 for noiseless ReLU networks resembles the phase diagram for linear transfer with deep linear networks in the noiseless setting with $\sigma=0$ (Appendix G Fig.7. 8). Overall, this demonstrates that we are able to predict the phase boundary between positive and negative transfer in the ReLU case, using our conceptual understanding in the deep linear case.
@@ 5 Conclusion
In this paper, we highlight the importance of thinking about transfer learning in the context of the feature space of the pretrained model. We rigorously identify the number of data points necessary for transfer learning to outperform scratch training as a function of feature space overlap in deep linear networks. We also demonstrate that our understanding of linear transfer carries over to shallow nonlinear networks as well. One of our primary findings is that transferability is inherited from the learned features of the pretraining task. In the rich training regime, this can lead to an inability for the pretrained model to transfer to tasks outside the source feature space. On the other hand, a model trained in the lazy regime is unlikely to outperfrom scratch training, since features are not updated in this limit. This suggests that models trained somewhere along the lazy-to-rich hierarchy may be more flexible in their transfer capabilities. In Appendix G Fig. 9 we generate a sweep of nonlinear models trained with varying degrees of feature learning on the source task and show that we can eliminate negative transfer if the pretrained model lies optimally between the lazy and rich regimes. These experiments demonstrate that regularizing pretrained models to avoid feature sparsification in the source task is a promising direction for improving transfer learning capabilities. With this work, we hope to advance the idea that transfer learning performance should be understood through the learned feature space of the pretrained model and not as a property of the dataset alone.
Impact Statement
This work aims to understand the theoretical aspects of transfer learning, and advances the idea that the feature space of the pretrained model determines the capability of the model to transfer to downstream tasks. This theoretical understanding could be used to create data and compute efficient pretraining procedures that maximize transferability to downstream tasks of interest.
Appendix A Dataset similarity is not predictive of transfer efficiency
The common wisdom in transfer learning is that related tasks should transfer effectively to one another. While it may seem that closeness of task distributions should correlate with transfer performance, we show that this is not necessarily the case. In particular, we select a member of each family and prove that, within our model, one can achieve positive transfer ($\mathcal{T}>0$) with distributions that are arbitrarily far apart. Two functions representable with the same features can be “far apart”. We formalize this notion with the following theorem.
Assumption A.1.
Theorem A.2.
Assume A.1. Then for any $f\in\Phi$, and any $\delta>0$ there exists $g\in\Phi$ such that
$\gamma_{\beta}(p_{f},p_{g})\geq\delta$
$D_{\mathrm{KL}}(p_{f}\|p_{g})\geq\delta$
where $D_{\mathrm{KL}}(p_{f}\|p_{g})$ is the Kullback Leibler divergence.
We prove this theorem in Appendix D.1. We note that this theorem also holds for any IPM over a function class that is larger than the class of Bounded Lipschitz functions. In particular, the theorem holds for the Monge-Kantorovich ($W_{1}$) metric, since any function that satisfies $\lVert f\rVert_{\mathrm{BL}}\leq 1$ also satisfies $\lVert f\rVert_{L}\leq 1$.
Theorem A.2 demonstrates that for a given source distribution, one can always find a target distribution generated from the same feature space that is arbitrarily distant with respect to these metrics, perhaps creating the illusion that transfer is likely to fail. However, even when the distance is large, if the source and target functions lie in the same feature space and pretraining creates a basis for this space, transfer to the target task will be positive, since only the output weights need to be relearned in the target task. We show this is indeed the case for deep linear networks in the following section.
Appendix B Initialization assumption
Following (Yun et al., 2021) we place the following constraint on the initialization for some $\lambda>0$.